# ANPrompt: Anti-noise Prompt Tuning for Vision-Language Models

### v2

@@ Introduction
1. Weak Noise Frozen Text Feature Generation: ANPrompt samples two intra-class descriptions: a main text that encodes core semantics, and a noise text that introduces subtle phrasing variations (e.g., attribute reordering). These two texts are encoded by the frozen CLIP text encoder into $f_{m}$ and $f_{n}$ respectively, and fused as follows:
$f_{w}=f_{m}+\alpha f_{n}$ (2)
@@
where $\alpha$ regulates the perturbation strength. Since the Weak Noise Frozen Text Feature $f_{w}$ incorporates contextual information from different sentences, it inevitably captures contextual cues that lean toward semantically adjacent yet incorrect classes. Consequently, weak noise text features inherently encode weak semantic perturbations information.
@@
2. Anti-noise Prompt Generation: To map weak semantic perturbation information into the prompt space, ANPrompt performs K-means clustering on all $f_{w}$ to extract representative semantic centers, which are referred to as noise prompts. These noise prompts are added to the learnable prompt tokens $P_{c}$ to form intermediate anti-noise prompts $P_{a}$. These intermediate prompts are then fed into two separate fully connected layers to generate modality-specific Anti-noise Prompts, which is injected into the deeper layers of both encoders.
3. Feature Encoding. After prompt injection, we extract three types of features: the Prompted Image Feature $f_{v}$, encoded from the final image class token; the Prompted Text Feature $f_{t}$, derived from the final [EOS] token; and the Noise-Resistant Visual Prompt Prototype (NRVPP) $f_{N}$, obtained by mean-pooling the visual Anti-noise Prompts after encoding, which captures weak semantic perturbations information.
4. Logit Computation. We compute four sets of logits: the Alignment Logits $\ell_{A}=f_{v}\cdot f_{t}^{\top}$, which measure the similarity between the prompted image and text features; the Robustness Logits $\ell_{R}=f_{v}\cdot f_{w}^{\top}$, which assess the model’s robustness against weak semantic noise in text; the Weak Noise Logits $\ell_{W}=f_{N}\cdot f_{w}^{\top}$, which reflect the alignment between noise-resistant visual features and weakly perturbed text features; and the Anti-noise Logits $\ell_{\text{final}}=\theta\ell_{A}+(1-\theta)\ell_{R}$, which integrate alignment and robustness for final prediction.
To stabilize training under weak semantic perturbations, we apply $\mathcal{L}_{\text{WA}}$ between $\ell_{W}$ and $\ell_{A}$ for alignment consistency under weak semantic perturbations, $\mathcal{L}_{\text{CE}}$ on $\ell_{\text{final}}$ for task supervision, and $\mathcal{L}_{\text{sim}}$ between frozen and prompted features to enhance semantic consistency.
The final loss is:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{sim}}+\gamma\mathcal{L}_{\text{WA}}$ (3)
where $\mathcal{L}_{\text{CE}}$ denotes a cross-entropy loss, $\mathcal{L}_{\text{sim}}$ represents a cosine loss, and $\mathcal{L}_{\text{WA}}$ is defined as follows. $\lambda$ serves as the weight of $\mathcal{L}_{\text{sim}}$, while $\gamma$ is a variance-adaptive weight used to regulate the strength of weak semantic perturbations.
@@
Weak Noise Alignment Loss
@@
To enforce alignment consistency under weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss ($\mathcal{L}_{\text{WA}}$) that minimizes the divergence between $\ell_{W}$ and $\ell_{A}$. Since KL divergence measures the discrepancy between distributions, minimizing it encourages the perturbed prediction $\ell_{W}$ to stay close to the original alignment $\ell_{A}$, penalizing semantic shifts and improving robustness while preserving alignment structure.
To compute $\mathcal{L}_{\text{WA}}$, we first derive $\ell_{A}$ from the prompted image and text features. Then, to obtain $\ell_{W}$, we compute the noise-resistant visual prototype $f_{N}$ by averaging the visual prompt token features $f_{v}^{o}\in\mathbb{R}^{N\times C}$:
$f_{N}=\frac{1}{N}\sum_{i=1}^{N}f_{v}^{o(i)}$ (4)
This averaging suppresses token-level variations and mitigates visual semantic distraction, especially under anti-prompts with weak semantic noise, where individual tokens may attend to irrelevant regions such as backgrounds or co-occurring objects. The weak noise logits are then computed as:
$\ell_{W}=f_{N}\cdot f_{w}^{\top}$ (5)
Finally, the weak alignment loss is defined as:
$\mathcal{L}_{\text{WA}}=\gamma\cdot\text{KL}(\text{softmax}(\ell_{R})\|\text{softmax}(\ell_{W}))$ (6)
with the variance-adaptive weight $\gamma$ given by:
$\gamma=\frac{1}{\text{std}(\ell_{R})\cdot|\ell_{R}|+\epsilon_{0}}$ (7)
@@
where $\text{std}(\ell_{R})$ denotes the standard deviation over all elements in $\ell_{R}$, $|\ell_{R}|$ represents the total number of elements (i.e., numel() of logits), and $\epsilon_{0}$ is a small constant to prevent division by zero. This global scaling dynamically adjusts the loss weight based on the dispersion and size of the logits, stabilizing optimization under varied prediction confidence levels.
Weak Semantic Perturbations Prompt Injection Mechanism
We introduce a prompt injection mechanism based on weak semantic perturbations. Specifically, we first cluster the weak noise frozen text features $f_{w}\in\mathbb{R}^{N\times C}$ into $K$ noise prompts $P_{w}\in\mathbb{R}^{k\times C}$ using K-Means, where $k$ matches the number of learnable prompt tokens $P_{c}\in\mathbb{R}^{k\times C}$ to ensure one-to-one integration.
These noise prompts are added to the learnable tokens to form the anti-noise prompt:
$P_{a}=P_{c}+\epsilon P_{w}$ (8)
@@
where $\epsilon$ controls the strength of injected weak perturbations. The anti-noise prompt $P_{a}$ is then projected into modality-specific spaces via a mapping layer.
Finally, the projected prompts $P_{a}^{\text{image}}$ and $P_{a}^{\text{text}}$ are injected into the visual and textual input sequences:
$\{P_{a}^{\text{image}},e_{\text{cls}},e_{1},e_{2},\dots,e_{M}\}$ (9)
$\{t_{\text{SOS}},P_{a}^{\text{text}},t_{1},t_{2},\dots,t_{L},c_{k},t_{\text{EOS}}\}$ (10)
@@
This injection enriches the prompt space with weakly perturbed semantic cues, promoting robustness against subtle noise while maintaining structural alignment with the original prompt format.
@@
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bold. We can see that our method achieves the best performance compared with other state-of-the-art methods.
Method
Ref.
Average ImageNet Caltech101 OxfordPets
@@
Method
Ref.
StandfordCars Flowers102 Food101 FGVCAircraft
@@
Table 2: Results from cross-dataset evaluation. ANprompt provides the highest average accuracy.
@@
Experiments
Evaluation Task
Base-to-Novel Generalization
To assess the zero-shot generalization capability of ANprompt, we divide each dataset into base and novel categories. The model is trained on the base categories using a few-shot setting, and its performance is evaluated on both the base and novel categories without additional training.
Cross-Dataset Evaluation
To examine the transferability of the model across datasets, we conduct experiments where the model is trained on a few-shot subset of the ImageNet dataset and directly evaluated on other datasets, without any further adaptation or fine-tuning.
Domain Generalization
To test the model’s resilience to domain shift, we use the ImageNet-trained model and evaluate it on four distinct ImageNet-derived datasets that introduce various domain differences, allowing us to measure performance under out-of-distribution scenarios.
Evaluation Settings
Datasets
@@
In the tasks of Base-to-New Generalization and Cross-Dataset Evaluation, we followed the experimental setup of CoOp (Zhou et al. 2022b), and evaluated on 11 public datasets, including ImageNet, Caltech (Fei-Fei, Fergus, and Perona 2004), OxfordPets (Parkhi et al. 2012), Flowers (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Gool 2014), StanfordCars (Krause et al. 2013), FGVCAircraft (Maji et al. 2013), EuroSAT (Helber et al. 2019), UCF101 (Kay et al. 2017), DTD (Cimpoi et al. 2014) and SUN397 (Xiao et al. 2010). We also use ImageNet (Deng et al. 2009) as the source dataset, and use its four variants ImageNetV2 (Recht et al. 2019), ImageNet-Sketch (Wang et al. 2019), ImageNet-A (Hendrycks et al. 2021b) and ImageNet-R (Hendrycks et al. 2021a) as the target dataset to evaluate the domain generalization task.
Implementation Details
@@
We set the learning rate to 0.001 and adopt the Adam optimizer to train ANPrompt. The model is trained for 10 epochs with a batch size of 4, using ViT-B/16 as the vision backbone. For the LLM cache, we incorporate prompts from CoPrompt (Roy and Etemad 2024) and HPT (Wang et al. 2024). And we set the $\theta$ and $\alpha$ with 0.7 and 0.001. All experiments are conducted on a single NVIDIA RTX 3090 Ti GPU. Reported results are averaged over three random seeds.
Experimental Performance
We compare the results of our method on three evaluation tasks with other advanced prompt learning methods, including zero-shot CLIP and some prompt learning methods, such as CoOp (Zhou et al. 2022b), CoCoOp (Zhou et al. 2022a), MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b), CoPrompt (Roy and Etemad 2024), TextRefiner (Xie et al. 2025) and MMRL (Guo and Gu 2025). We also added some targeted ablation experiments to prove the effectiveness of the proposed ANPrompt method.
Base-to-New Generalization
@@
Table 1 shows the performance of ANPrompt and other comparison methods on B2N generalization tasks. Compared with the most advanced MMRL method, ANPrompt improves the base class by 0.44% and the new class by 0.43%. Combining the base class and the new class, ANPrompt achieves a performance improvement of 0.44% on the harmonic average. It is worth noting that compared with MMRL (Guo and Gu 2025), ANPrompt achieves significant performance improvement on EuroSAT dataset: 6.03% on new class and 3.31% on harmonic average. By transforming weak semantic noise into constructive training signals, ANPrompt achieves performance improvement by fusing complementary semantic perturbation sentences.
Cross-Dataset Evaluation
@@
Table 2 shows the performance comparison between our ANPrompt and existing methods in cross-dataset evaluation. On the ImageNet source dataset, ANPrompt shows comparable performance to other competitive methods. On the target domain dataset, ANPrompt achieved the best performance on six of the ten datasets, showing excellent generalization ability, especially on fine-grained datasets, such as FGVCAircraft and EuroSAT. Specifically, with an average of 10 datasets, ANPrompt achieved the highest accuracy rate of 67.14%. Meanwhile, compared with the suboptimal method CoPrompt, ANPrompt achieved 0.14% performance improvement. Different from the existing methods, we use the anti-noise visual prompt prototype to align the weak noise frozen text features to generate soft, context-aware supervision signals, thus improving cross-domain performance.
Table 3: Results on domain generalization task.
@@
Domain Generalization
We compare the performance of our ANPrompt with that of six previous methods on cross-domain data sets. As shown in Table 3, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability. Specifically, ANPrompt performs best on Imagenet-V2 dataset, and it also shows considerable performance on other target domain datasets.
Ablation Experiments
Componment Ablation
We conduct an ablation study to evaluate the contribution of each component in ANPrompt, including the Weak noise Frozen Text feature (TextNoise), Weak Alignment Loss (WALoss), and Anti-noise Prompt injection. As shown in Table 4, introducing WALoss or Anti-Prompt individually yields consistent improvements over the baseline, with gains of +0.06 and +0.19 respectively. Notably, the addition of TextNoise leads to further gains when combined with either WALoss (+0.34) or Anti-Prompt (+0.40). When all three components are enabled, ANPrompt achieves the highest harmonic mean (HM) of 81.70, surpassing the baseline by +0.79. These results verify the complementary nature of each module and their collective effectiveness in enhancing robustness under noise.
Table 4: The Effectiveness of different components in ANPrompt. The TextNoise stands for the Weak noise Frozen Text feature.

### v3

@@ III-B1 Weak Noise Frozen Text Feature Generation
ANPrompt samples two intra-class descriptions: a main text that encodes the core semantics and a noise text that introduces subtle phrasing variations (e.g., attribute reordering). These two texts are encoded by the frozen CLIP text encoder into features $f_{m}$ and $f_{n}$, which are then fused as:
$f_{w}=f_{m}+\alpha f_{n},$ (2)
@@
where $\alpha$ is a scaling factor that controls the strength of the perturbation. The Weak Noise Frozen Text Feature $f_{w}$ combines contextual information from both the main and noise texts, capturing weak semantic perturbations that subtly shift semantics toward adjacent, but incorrect, classes.
The main and noise texts are randomly sampled from a diverse intra-class language cache. This approach effectively captures natural linguistic variations, approximating real-world weak perturbations. Future work may explore more controlled alternatives, such as paraphrasing or synonym replacement, to generate more fine-grained noise variations.
III-B2 Anti-noise Prompt Construction
@@
To map weak semantic perturbation information into the prompt space, ANPrompt applies K-means clustering on all $f_{w}$ to extract representative semantic centers, denoted as noise prompts. These noise prompts are then added to the learnable prompt tokens $P_{c}$ to form intermediate anti-noise prompts $P_{a}$. Two modality-specific fully connected layers then transform $P_{a}$ into Anti-noise Prompts for the vision and text encoders, which are injected into their deeper layers.
III-B3 Feature Encoding
After prompt injection, three types of features are obtained:
  • Prompted Image Feature $f_{v}$, encoded from the final image class token;
  • Prompted Text Feature $f_{t}$, derived from the [EOS] token;
  • Noise-Resistant Visual Prompt Prototype (NRVPP) $f_{N}$, obtained by mean-pooling the visual Anti-noise Prompts. This step helps capture weak semantic perturbations while suppressing high-frequency token-level variations, ensuring the preservation of meaningful visual information across perturbations.
III-B4 Logit Computation
We compute four sets of logits:
  • Alignment Logits $\ell_{A}=f_{v}\cdot f_{t}^{\top}$, measuring the similarity between prompted image and text features;
  • Robustness Logits $\ell_{R}=f_{v}\cdot f_{w}^{\top}$, assessing the model’s robustness to weak semantic noise in the text;
  • Weak Noise Logits $\ell_{W}=f_{N}\cdot f_{w}^{\top}$, reflecting the alignment between noise-resistant visual features and weakly perturbed text features;
  • Anti-noise Logits $\ell_{\text{final}}=\theta\ell_{A}+(1-\theta)\ell_{R}$, integrating alignment and robustness for the final prediction.
After computing these four logits, we can compute the training losses: cross-entropy loss ($\mathcal{L}_{\text{CE}}$), weak alignment loss ($\mathcal{L}_{\text{WA}}$), and cosine similarity loss ($\mathcal{L}_{\text{sim}}$). Specifically:
  • The cross-entropy loss ($\mathcal{L}_{\text{CE}}$) is computed using the Alignment Logits $\ell_{A}$.
@@ III-B4 Logit Computation
  • The weak alignment loss ($\mathcal{L}_{\text{WA}}$) is computed based on the Weak Noise Logits $\ell_{W}$ and the Alignment Logits $\ell_{A}$.
  • The cosine similarity loss ($\mathcal{L}_{\text{sim}}$) is calculated between the image and text embeddings to preserve semantic consistency across modalities.
The overall training loss is:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{sim}}+\gamma\mathcal{L}_{\text{WA}},$ (3)
where $\lambda$ controls the contribution of $\mathcal{L}_{\text{sim}}$, and $\gamma$ adaptively controls the effect of $\mathcal{L}_{\text{WA}}$.
where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\text{sim}}$ is the cosine similarity loss, and $\mathcal{L}_{\text{WA}}$ is the weak alignment loss. Here, $\mathcal{L}_{\text{sim}}$ is defined over both image and text embeddings to preserve semantic consistency across modalities. In addition, $K$ explicitly denotes the number of learnable prompt tokens, ensuring one-to-one integration with the clustered noise prompts.
III-C Weak Alignment Loss
@@
To enforce consistency under weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss ($\mathcal{L}_{\text{WA}}$) that minimizes the divergence between $\ell_{W}$ and $\ell_{A}$. Minimizing this divergence encourages predictions under perturbation to remain close to the clean alignment, penalizing semantic shifts while preserving structural consistency.
To compute $\ell_{W}$, the NRVPP $f_{N}$ is obtained by averaging visual prompt token features $f_{v}^{o}\in\mathbb{R}^{N\times C}$:
$f_{N}=\frac{1}{N}\sum_{i=1}^{N}f_{v}^{o(i)},$ (4)
which suppresses token-level variations and mitigates visual semantic distraction caused by weak noise. The weak noise logits are then:
$\ell_{W}=f_{N}\cdot f_{w}^{\top}.$ (5)
The Weak Alignment Loss is defined as:
$\mathcal{L}_{\text{WA}}=\text{KL}(\text{softmax}(\ell_{A})\,\|\,\text{softmax}(\ell_{W})),$ (6)
We use $\gamma$ to control the variance-adaptive weight:
$\gamma=\frac{1}{\text{std}(\ell_{R})\cdot|\ell_{R}|+\epsilon_{0}},$ (7)
@@
where $\text{std}(\ell_{R})$ is the standard deviation of $\ell_{R}$, $|\ell_{R}|$ is its number of elements, and $\epsilon_{0}$ is a small constant for numerical stability. This dynamic scaling adjusts the loss weight based on prediction dispersion, stabilizing optimization across varied confidence levels.
III-D Weak Semantic Perturbation Prompt Injection
Finally, we formalize the injection mechFanism. The weak noise features $f_{w}\in\mathbb{R}^{N\times C}$ are clustered into $K$ noise prompts $P_{w}\in\mathbb{R}^{K\times C}$ using K-means, with $K$ matching the number of learnable prompt tokens $P_{c}\in\mathbb{R}^{K\times C}$. The anti-noise prompts are then defined as:
$P_{a}=P_{c}+\epsilon P_{w},$ (8)
@@ III-D Weak Semantic Perturbation Prompt Injection
where $\epsilon$ controls the perturbation strength. $P_{a}$ is projected into modality-specific spaces to obtain $P_{a}^{\text{image}}$ and $P_{a}^{\text{text}}$, which are injected into the input sequences:
$\{P_{a}^{\text{image}},e_{\text{cls}},e_{1},\dots,e_{M}\},$ (9)
$\{t_{\text{SOS}},P_{a}^{\text{text}},t_{1},\dots,t_{L},c_{k},t_{\text{EOS}}\}.$
@@
This injection enriches the prompt space with weakly perturbed cues, promoting robustness against subtle noise while maintaining structural alignment with the original prompt format.
@@
TABLE I: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bold. We can see that our method achieves the best performance compared with other state-of-the-art methods.
Method Ref. Average ImageNet Caltech101 OxfordPets
@@
Method Ref. StandfordCars Flowers102 Food101 FGVCAircraft
@@
TABLE II: Results from cross-dataset evaluation. ANPrompt provides the highest average accuracy.
@@
IV EXPERIMENT
IV-A Dataset
@@
In the tasks of Base-to-New Generalization and Cross-Dataset Evaluation, we followed the experimental setup of CoOp coop , and evaluated on 11 public datasets, including ImageNet, Caltech caltech101 , OxfordPets oxfordpets , Flowers Flowers102 , Food101 food101 , StanfordCars StanfordCars , FGVCAircraft fgvcaircraft , EuroSAT Eurosat , UCF101 ucf101 , DTD dtd and SUN397 sun . We also use ImageNet imagenet as the source dataset, and use its four variants ImageNetV2 imagenetv2 , ImageNet-Sketch imagenets , ImageNet-A imageneta and ImageNet-R imagenetr as the target dataset to evaluate the domain generalization task.
IV-B Implementation Details
@@
We set the learning rate to 0.001 and adopt the Adam optimizer to train ANPrompt. The model is trained for 10 epochs with a batch size of 4, using ViT-B/16 as the vision backbone. For the LLM cache, we incorporate prompts from CoPrompt coprompt and HPT hpt . And we set the $\theta$ and $\alpha$ with 0.7 and 0.001. We inject the prompt into the layer 6 to 12. All experiments are conducted on a single NVIDIA RTX 3090 Ti GPU. Reported results are averaged over three random seeds.
IV-C Base-to-New Generalization
In this experiment, we evaluate ANPrompt and benchmark its performance against several existing methods, such as CLIP clip , CoOp coop , CoCoOp cocoop , MaPLe maple , PromptSRC promptsrc , CoPrompt coprompt , TextRefiner textrefiner , MMRL mmrl , SurPL-G surpl , LwEIB lweib and BIP-D bip .
@@ IV-C Base-to-New Generalization
TABLE I shows the performance of ANPrompt and other comparison methods on B2N generalization tasks. Compared with the most advanced LwEIB method, ANPrompt improves the base class by 1.7% and the harmonic average by 0.49%. It is worth noting that compared with BIP-D, ANPrompt achieves significant performance improvement on EuroSAT dataset: 9.33% on new class and 5.26% on harmonic average. By transforming weak semantic noise into constructive training signals, ANPrompt achieves performance improvement by fusing complementary semantic perturbation sentences.
IV-D Cross-Dataset Generalization
@@
TABLE II shows the performance comparison between our ANPrompt and existing methods in cross-dataset evaluation. On the ImageNet source dataset, ANPrompt shows comparable performance to other competitive methods. On the target domain dataset, ANPrompt achieved the best performance on four of the ten datasets, showing excellent generalization ability, especially on fine-grained datasets, such as FGVCAircraft and EuroSAT. Specifically, with an average of 10 datasets, ANPrompt achieved the highest accuracy rate of 67.14%. Meanwhile, compared with the suboptimal method CoPrompt, ANPrompt achieved 0.14% performance improvement. Different from the existing methods, we use the anti-noise visual prompt prototype to align the weak noise frozen text features to generate soft, context-aware supervision signals, thus improving cross-domain performance.
TABLE III: Results on domain generalization task.
@@
IV-E Domain Generalization
We compare the performance of our ANPrompt with that of nine previous methods on cross-domain data sets. As shown in TABLE III, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability. Specifically, ANPrompt performs best on Imagenet-V2 dataset, and it also shows considerable performance on other target domain datasets.
IV-F Few-shot Learning
We evaluate the discriminating ability of the model under limited supervision by changing the number of shots in each category (K = 1, 2, 4, 8, 16). The evaluation includes five baseline methods and ANPrompt proposed by us, such as CLIP, CoOp, CoCoOp, Prograd and Maple. As shown in Figure 6, ANPrompt is always superior to the comparison method in all lens settings, which shows the robustness of ANPrompt. Specifically, compared with the suboptimal method Maple, ANPrompt achieved an average performance improvement of 2.76%, 2.29%, 1.99%, 0.92% and 0.65% on 1, 2, 4, 8 and 16 shots on 11 data sets, respectively. In terms of a single data set, ANPrompt showed the best accuracy on 8 of 11 datasets. For example, in 16shots, ANPrompt achieved 74.27% value on DTD dataset, which was 2.94% higher than Maple. Compared with Maple, UCF101 is improved by 1.37%. In addition, ANPrompt’s performance on other datasets is still competitive. This shows that ANPrompt effectively enhances the robustness under weak semantic disturbance by injecting weak semantic disturbance into the alignment process and prompt representation.
Figure 6: Few-shot classification accuracy across multiple datasets under varying shots. ANPrompt consistently outperforms existing methods.
We compare the performance of our ANPrompt with that of six previous methods on cross-domain data sets. As shown in TABLE III, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability. Specifically, ANPrompt performs best on Imagenet-V2 dataset, and it also shows considerable performance on other target domain datasets.
IV-G Ablation Study
IV-G1 Component ablation
We conduct an ablation study to evaluate the contribution of each component in ANPrompt, including the Weak noise Frozen Text feature (TextNoise), Weak Alignment Loss (WALoss), and Anti-noise Prompt injection. As shown in Table IV, introducing WALoss or Anti-Prompt individually yields consistent improvements over the baseline, with gains of +0.06 and +0.19 respectively. Notably, the addition of TextNoise leads to further gains when combined with either WALoss (+0.34) or Anti-Prompt (+0.40). When all three components are enabled, ANPrompt achieves the highest harmonic mean (HM) of 81.70, surpassing the baseline by +0.79. These results verify the complementary nature of each module and their collective effectiveness in enhancing robustness under noise.
TABLE IV: The Effectiveness of different components in ANPrompt. The TextNoise stands for the Weak noise Frozen Text feature.